Good evening how would I Solve 2-3(4-x)=x-(3-x)
Answer by usyim88hk(157)

2-3(4-x) = x-(3-x)
First multiply 3 and (4-x)
2-12+3x = x-(3-x)
Then subtract 12 from 2
-10+3x = x-(3-x)
Now, subtract (3-x) from x
-10+3x = x-3+x
-10+3x = 2x-3
Now add three to both sides
-7+3x = 2x
subtract 3x from both sides
-7 = -x
Divide -1 from both sides
7 = x